# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

### v3

@@ 3 Fairness-Aware Adversarial Missing Data Imputation for Graphs
We introduce a graph learning framework, which we call BFtS (Better Fair than Sorry), that allows fair GNN training in a setting with missing sensitive data. BFtS applies a 3-player adversarial learning scheme, where the fair GNN is trained jointly with two adversaries—one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness. As a result, fairness assessments by the GNN are based on the worst-case imputed values generated by the adversarial imputation.
@@ 3.1 Proposed Model (BFtS)
Our goal is to learn a fair and accurate GNN $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ for node classification with missing sensitive values. The proposed solution is model-agnostic and based on adversarial learning. It formulates the imputation model as a second adversary of the GNN.
Figure 3 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{I}$, a node classification GNN $f_{C}$, and a sensitive attribute Deep Neural Net (DNN) prediction $f_{A}$. $f_{C}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{A}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{C}$ to assess the bias. More specifically, it $f_{C}$ is assumed to be biased if the adversary $f_{A}$ can accurately predict the sensitive attribute information from the representations of $f_{C}$. To train $f_{C}$, we need to know the values of the missing sensitive attributes. $f_{I}$ predicts the missing sensitive attributes. It takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and generates the missing sensitive attributes $\hat{si}$. $f_{I}$ aims to generate sensitive values that minimize the fairness of $f_{C}$. Therefore, it works as a second adversary to the model.
We will formalize the architecture of the players, define their loss functions, and describe the training procedure for our approach in the next section.
@@ 3.1.1 Player architectures
This section describes the architecture of each player in BFtS.
@@ 3.1.2 Loss Functions
Here, we describe the loss functions applied in our approach.
@@
Let $\hat{y}=f_{C}(X,G)$ be the predictions of $f_{C}$. We apply the cross-entropy loss to $f_{C}$’s predictions as follows:
@@
Because sensitive attributes tend to be imbalanced in real datasets, $f_{I}$ applies the Label-Distribution-Aware Margin (LDAM) loss Cao et al. (2019) for classification. Let $\hat{si}=f_{I}(X,G)$ be the one-hot encoded predictions and $\hat{si}^{s}_{v}$ be the output of node $v$ for $s_{v}$. Then, the LDAM loss of $f_{I}$ is defined as:
@@
where $\Delta^{j}=C/n_{j}^{\frac{1}{4}}$ and $j\in\{0,1\}$.
Here $C$ is a constant independent of the sensitive attribute $s$, and $n_{j}$ is the number of samples that belong to $s=j$. The LDAM loss is a weighted version of the negative log-likelihood loss. Intuitively, since $\Delta^{j}$ is larger for smaller values of $n_{j}$, it ensures a higher margin for the smaller classes.
Furthermore, BFtS can generate fair outputs when very little or no sensitive information is provided. When no or very little sensitive information is available, ($|\mathcal{V}_{S}|\to 0$), $f_{I}$ can impute the sensitive values from the ground truth label $y$ with LDAM loss. We can replace $\mathcal{V}_{S}$ with $\mathcal{V}_{L}$ and $s$ with $y$ for each node $v\in\mathcal{V}_{L}$ in Eqtn 2. The primary argument is that the worst-case fairness model generally assigns a more desired outcome to the non-protected group and a less desired outcome to the protected group. Therefore, the nodes predicted as the minority class will fall into the protected group, and vice versa. Compared to the model $f_{I}$ trained with LDAM loss with limited sensitive information of $\mathcal{V}_{S}$; without any sensitive information, $f_{I}$ trained with LDAM loss learns more bias-amplified samples with labels $y$ of $V_{L}$. Depending on the amount of sensitive information available, the user can choose any of these two methods to learn the imputation.
@@
$\mathcal{L}_{A}=$ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]$
$+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))]$ (3)
@@ 3.1.3 Learning the parameters of BFtS
Let $\theta_{C},\theta_{A}$, and $\theta_{I}$ be the parameters of models $f_{C}$, $f_{A}$, and $f_{I}$, respectively. We propose learning these parameters jointly via a 3-player adversarial learning scheme. Our goal is for $f_{I}$ to learn how to impute the missing sensitive values that represent the worst-fairness scenario for the node representations learned by $f_{C}$, which depend on the sensitive attribute predictions produced by $f_{A}$. This will induce $f_{C}$ to be both fair and accurate. Thus, $\theta_{C}$ are optimized as follows:
@@
The parameters of the sensitive attribute predictor $\theta_{A}$ only influence the $\mathcal{L_{A}}$ loss. $\theta_{A}$ are learned by maximizing $\mathcal{L}_{A}$:
@@
$\begin{split}\min_{\theta_{C}}\max_{\theta_{I},\theta_{A}}~{}&\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]\\ &+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))].\end{split}$ (7)
3.2 Training Algorithm

### v4

@@ 4 Fairness-Aware Adversarial Missing Data Imputation for Graphs
We introduce a framework called BFtS (Better Fair than Sorry) that allows fair GNN training with adversarial missing sensitive data. BFtS applies a 3-player adversarial learning scheme, where the fair GNN is trained jointly with two adversaries—one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness. As a result, fairness assessments by the GNN are based on the worst-case imputed values from the adversarial imputation.
@@ 4.1 Proposed Model (BFtS)
Figure 4: 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{C}$ generates node representations by minimizing the classification loss $\mathcal{L}_{C}$ (Eqn. 1) and the maximizing sensitive attribute prediction loss $\mathcal{L}_{A}$ (Eqn. 3). $f_{A}$ predicts sensitive attributes using representations from $f_{C}$ by minimizing $\mathcal{L}_{A}$. $f_{I}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{I}$ (Eqn. 2) and maximizing $\mathcal{L}_{A}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are predictions from $f_{C}$, $f_{I}$ and $f_{A}$, respectively.
Our goal is to learn a fair and accurate GNN $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ for node classification with missing sensitive values. The proposed solution is model-agnostic and based on adversarial learning. It formulates the imputation model as a second adversary of the GNN. Figure 4 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{I}$, a node classification GNN $f_{C}$, and a sensitive attribute Deep Neural Net (DNN) prediction $f_{A}$. $f_{C}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{A}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{C}$ to assess the bias. More specifically, it $f_{C}$ is assumed to be biased if the adversary $f_{A}$ can accurately predict the sensitive attribute information from the representations of $f_{C}$. To train $f_{C}$, we need to know the values of the missing sensitive attributes. The model $f_{I}$ predicts the missing sensitive attributes by taking $\mathcal{X}$ and $\mathcal{G}$ as inputs and generating the missing sensitive attributes $\hat{si}$. The goal of $f_{I}$ is to generate sensitive values that minimize the fairness of $f_{C}$ and, thus, it works as a second adversary to $f_{C}$.
@@ 4.1.2 Loss Functions
We apply the cross-entropy loss to learn $f_{C}$ as follows:
@@
Because sensitive attributes tend to be imbalanced, $f_{I}$ applies the Label-Distribution-Aware Margin (LDAM) loss ldam . Let $\hat{si}=f_{I}(X,G)$ be the one-hot encoded predictions and $\hat{si}^{s}_{v}$ be the output of node $v$ for $s_{v}$. The LDAM loss of $f_{I}$ is defined as:
@@
where $\Delta^{j}=C/n_{j}^{\frac{1}{4}}$ and $j\in\{0,1\}$, $C$ is a constant independent of the sensitive attribute $s$, and $n_{j}$ is the number of samples that belong to $s=j$.
The LDAM loss is a weighted version of the negative log-likelihood loss. Intuitively, since $\Delta^{j}$ is larger for smaller values of $n_{j}$, it ensures a higher margin for the smaller classes.
BFtS with no sensitive information: BFtS can generate fair outputs when very little or no sensitive information is provided by letting $f_{I}$ impute the sensitive values from the ground truth training labels $y$ using the LDAM loss. We replace $\mathcal{V}_{S}$ with $\mathcal{V}_{L}$ and $s$ with $y$ for each node $v\in\mathcal{V}_{L}$ in Eqn 2. The reasoning is that the worst-case fairness model generally assigns a more desired outcome to the non-sensitive group and a less desired outcome to the sensitive one. Therefore, the nodes predicted as the minority class will fall into the sensitive group, and vice versa. This is consistent with the worst-case assumption of BFtS. RNF du2021fairness applies a similar solution with a different loss function.
@@
$\mathcal{L}_{A}=\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))]$ (3)
@@ 4.1.3 Learning the parameters of BFtS
Let $\theta_{C},\theta_{A}$, and $\theta_{I}$ be the parameters of $f_{C}$, $f_{A}$, and $f_{I}$, respectively. We learn these parameters via a 3-player adversarial learning scheme described in Figure 4. Parameters $\theta_{C}$ are optimized as follows:
@@
The parameters of the sensitive attribute predictor $\theta_{A}$ are learned by maximizing $\mathcal{L}_{A}$:
@@
$\begin{split}\min_{\theta_{C}}\max_{\theta_{I},\theta_{A}}\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))].\end{split}$ (7)
The training algorithm and time complexity analysis of BFtS are discussed in Appendix A.
4.2 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are provided in Appendix A.
Theorem 2.
$f_{I}$ generates imputations such that $\Delta DP$ in Sec 5) is maximized for a given classifier $f_{C}$ .
Theorem 3.
The maximal $\Delta DP$ of BFtS is minimal among all estimators of $s$.
The optimal GNN classifier $f^{*}_{C}$, will achieve demographic parity ($\Delta DP=0$ in Sec 5) for the worst-case imputation (minimum fairness) $\hat{s}$ generated by $f_{I}$.
Corollary 1.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is related to the convergence of adversarial learning. For FairGNN, if $s^{\prime}$ is inaccurate, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{A}$ in Eq. 4 will be constant. The interplay between the three players makes BFtS more robust to convergence issues because the objective minimizes the upper bound on the JS divergence. This reduces the probability that the divergence vanishes during training (see details in Appendix A).
5 Experiments and Results
We compare our approach for learning fair GNNs for node classification (BtFS) against the state-of-the-art in terms of accuracy and fairness using real and synthetic data. The size of $\mathcal{V}_{S}$ was set to $30\%$ of $|\mathcal{V}|$ unless stated otherwise. The details of the datasets, baselines, hyperparameters, additional experiments with graph attention network, and ablation studies are in Appendix A.
RNF BFtS
%AVPR($\uparrow$) %F1($\uparrow$) $\%\Delta DP$($\downarrow$) $\%\Delta EQ(\downarrow$) %AVPR($\uparrow$) % F1($\uparrow$) $\%\Delta DP$ ($\downarrow$) $\%\Delta EQ(\downarrow)$
BAIL 81.0±0.1 85.3±0.3 11.65±0.02 8.51±0.02 83.1±0.2 86.2±0.3 8.01±0.03 4.12±0.01
Credit 80.4±0.4 75.9±0.1 8.01±0.08 7.25±0.08 82.1±0.4 76.8±0.0 5.97±0.10 4.96±0.06
German 73.2±0.2 74.4±0.1 9.42±0.04 8.92±0.12 74.1±0.3 74.7±0.4 6.42±0.11 7.68±0.17
NBA 70.1±0.2 70.2±0.1 6.47±0.03 5.89±0.05 72.9±0.3 69.8±0.1 5.19±0.01 3.59±0.05
Pokek-z 73.1±0.4 71.2±0.1 6.18±0.05 6.29±0.07 73.4±0.2 73.2±0.3 5.10±0.02 3.20±0.06
Pokek-n 71.6±0.2 72.2±0.1 7.58±0.01 7.09±0.04 72.6±0.1 69.6±0.4 4.29±0.02 3.01±0.03
Table 1: AVPR, F1 score, $\%\Delta DP$, and $\%\Delta EQOP$ without any sensitive information of BFtS and RNF. Amongst the baselines, only RNF works without any sensitive information. BFtS outperforms RNF in terms of fairness and accuracy for all baselines.
Evaluation: For utility evaluation, we use average precision (AVPR) and F1 score. For measuring bias we use $\Delta DP$ and $\Delta EQOP$. Details of these metrics are in Appendix A.
5.1 Results and Analysis
6 Conclusion
We present BFtS, a 3-player adversarial learning framework for the imputation of adversarially missing sensitive attributes that produce challenging values for graph-based fairness. Our theoretical analysis showed that our approach is more robust to convergence challenges during adversarial training than existing alternatives. We evaluate BFtS using synthetic and real datasets. Our experimental results provide strong evidence that our approach achieves a better fairness $\times$ accuracy trade-off than the baselines.
Limitations and societal impact: We note the following as limitations of our work: (1) our experiments do not consider large-scale graph datasets for fairness as they are currently unavailable, and (2) our theoretical analysis shows that BFtS achieves demographic parity at equilibrium but not other fairness notions. In terms of social impact, our approach enables fair graph ML with missing data. However, one can leverage our method to predict private information.
Appendix A Technical Appendices and Supplementary Material
This supplementary material includes the following:
  • Details of Graph Neural Network (GNN)
  • Evaluation Metrics
  • Dataset Details
  • Baselines
  • Software and Hardware
  • Hyperparameter Setting
  • Proof of theorems and corollaries
  • Training algorithm
  • Complexity and Running time comparison
  • Imputation accuracy
  • Additional experiments
  • Ablation Study
Graph Neural Network
Two of the models (or players) learned by our model are GNNs. Without loss of generality, we will assume that they are message-passing neural networks [25], which we will simply refer to as GNNs. A GNN learns node representations $\mathbf{h}_{v}$ for each node $v$ that can be used to predict (categorical or continuous) values $y_{v}=f(\mathbf{h}_{v})$. These representations are learned via successive $\mathsf{AGGREGATE}$ and $\mathsf{COMBINE}$ operations defined as follows:
Evaluation Metrices
Utility Evaluation: As with many datasets in the fairness literature, our datasets are imbalanced. To evaluate their accuracy, we apply average precision (AVPR) and the F1 score.
Fairness Evaluation: In machine learning literature, fairness is often measured with bias. We apply two bias metrics that will be defined based on a sensitive attribute $s$. Let $s=1$ be the sensitive group and $s=0$ be the non-sensitive group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute [36]. We evaluate demographic parity as:
$\Delta DP(\hat{y},s)=|p(\hat{y}=1|s=1)-p(\hat{y}=1|s=0)|.$ (8)
Equality of Opportunity (EQOP) calculates the difference between true positive rates between sensitive and non-sensitive groups [28] as follows:
$\begin{split}\Delta EQOP(\hat{y},s,y)=|p(\hat{y}=1|s=1,y=1)-p(\hat{y}=1|s=0,y=1)|.\end{split}$ (9)
We can get corresponding fairness by subtracting $\Delta DP$ and $\Delta EQOP$ from $1$.
Dataset Details
Simulation: We simulate a synthetic graph consisting of $1000$ nodes based on the stochastic block model with blocks of sizes $600$ and $400$, for the majority and minority classes, respectively. We can have different edge probabilities within and across blocks and generate graphs with different assortativity coefficients. The sensitive attribute $s$ is simulated as $s\!\sim\!Bernoulli(p)$ if $y=1$ and $s\!\sim\!Bernoulli(1-p)$ otherwise. The probability $p=0.5$ generates a random assignment of the sensitive attribute, while $p$ close to $1$ or 0 indicates an extremely biased scenario. For our simulation, we use $p=0.7$ so that the dataset is unfair. Each node has 20 attributes, of which 8 are considered noisy. The noisy attributes are simulated as $Normal(0,1)$ and the remaining ones as $\gamma*y+Normal(0,1)$, where $\gamma$ is the strength of the signal in important features.
Bail [34]: Contains 18,876 nodes representing people who committed a criminal offense and were granted bail by US state courts between 1990 and 2009. Defendants are linked based on similarities in their demographics and criminal histories. The objective is to categorize defendants as likely to post bail vs. no bail. No-bail defendants are expected to be more likely to commit violent crimes if released. Race information is considered a sensitive attribute. Features include education, marital status, and profession.
Credit [74]: The dataset has $30,000$ nodes representing people each with $25$ features, including education, total overdue counts, and most recent payment amount. The objective is to forecast if a person will or won’t miss a credit card payment while considering age as a sensitive attribute. Nodes are connected based on payment patterns and spending similarities.
German [30]: The dataset consists of 1000 nodes and 20 features (e.g., credit amount, savings, profession) representing clients at a German bank. Labels are credit decisions for a person (good or bad credit). The sensitive attribute is gender. The goal is to predict who should receive credit.
NBA [14]: Dataset based on a Kaggle dataset about the National Basketball Association (US). It contains 403 players as nodes connected based on their Twitter activity. Attributes represent player statistics and nationality is considered the sensitive attribute. The goal is to predict the players’ salaries (above/under the median).
Pokec [63]: This dataset has two variations: Pokec-z and pokec-n. The pokec dataset is based on a popular social network in Slovakia. The sensitive attribute in pokec-z and pokec-n are two different regions of Slovakia. The goal is to predict the working field of users.
Baselines
Vanilla: The vanilla model is the basic GNN model trained with cross-entropy loss without any fairness intervention.
Debias [77]: Debias is an adversarial learning method that trains a classifier and an adversary simultaneously. The adversary is trained using the softmax output of the last layer of the classifier. It does not handle missing sensitive data. Hence, we only use $\mathcal{V}_{S}\in\mathcal{V}$ to calculate the adversarial loss.
FairGNN [14]: FairGNN first uses a GNN to estimate the missing sensitive attributes. It then trains another GNN-based classifier and a DNN adversary. The adversary helps to eliminate bias from the representations learned by the GNN classifier.
RNF [17]: RNF imputes the missing sensitive attributes by first training the model with generalized cross entropy [80]. It eliminates bias from the classification head. RNF uses samples with the same ground-truth label but distinct sensitive attributes and trains the classification head using their neutralized representations.
FairVGNN [69]: FairVGNN learns fair node representations via automatically identifying and masking sensitive-correlated features. It requires complete sensitive information, so we use the same imputation method as FairGNN for the missing values.
FairSIN [73]: FairSIN learns fair representations by adding additional fairness-facilitating features. It also requires complete sensitive information, so we use the same independent imputation method as FairGNN for the missing values. [73]:
Software and Hardware
  • Operating System: Linux (Red Hat Enterprise Linux 8.9 (Ootpa))
  • GPU: NVIDIA A40
  • Software: Python 3.8.10, torch 2.2.1, dgl==0.4.3
Hyperparameter settings
Each GNN has two layers and a dropout probability of $0.5$. In practice, some sensitive information is often available, unless otherwise stated. Therefore, we concentrate the majority of our experiments on the assumption that there is a limited amount of sensitive information available, and we generate the imputations using the available sensitive information in Equation 2. The $\mathcal{V}_{S}$ was set to $30\%$ for all experiments unless stated otherwise. $\mathcal{V}_{L}$ is $300,400,1400,100,500,500$, and $800$ for the Simulation, German, Credit, NBA, pokec-z, pokec-n and Bail datasets, respectively. We adjust the values of $\alpha$ and $\beta$ for each dataset using cross-validation based on the F1 score. For all baselines, we choose the respective regularizers using cross-validation based on the same metrics. Every model, except RNF, combines the training of several networks with various convergence rates. We stop training RNF based on early stopping to eliminate overfitting. We train the rest of the models for certain epochs and choose the model with the best AVPR. All the results shown in the following section are the average of $10$ runs. We use GCN as the GNN architecture for all experiments unless mentioned otherwise.
Proof of Theorems
Theorem 1
We will consider an adversarial active learning setting, where the goal of the adversary is to select $k$ labels (i.e., observed sensitive values) to minimize the bias $\mathcal{L}_{A}$ of the data imputed by $f_{I}$. More specifically, our proof will use the coverage model for active learning, where an unlabeled data point $x_{i}^{u}$ can be correctly predicted iff it is covered by at least one labeled point $x_{j}^{l}$, so the goal is for the labeled points $\{x_{1}^{l},x_{2}^{l},\ldots x_{k}^{l}\}$ to cover as many unlabeled data points $\{x_{1}^{u},x_{2}^{u},\ldots x_{n-k}^{u}\}$ as possible [75, 57]. Thus, in our setting, the adversary’s goal is to select labels to minimize the coverage of the following sets of nodes: (1) nodes in the sensitive group ($s_{v}=1$) and the negative class ($y=0$) and (2) nodes in the non-sensitive group ($s_{v}=0$) and the positive class ($y=1$). Without loss of generality, we will focus on the second set $(s_{v}=0,y=1)$ because it impacts demographic parity as a bias metric. This can be achieved by minimizing the coverage of such nodes. We provide a reduction from the minimum k-union problem.
Minimum k-union problem (MkU, [13]): Given a set of sets $\mathcal{S}={S_{1},S_{2},\ldots S_{q}}$, where each set $S_{i}\subseteq\mathcal{I}$ and $\mathcal{I}$ is a set of items. The problem consists of selecting $r$ sets $S_{1},S_{2},\ldots S_{r}$ from $\mathcal{S}$ to minimize the coverage $S_{1}\cap S_{2}\cap\ldots S_{r}$.
Under our coverage model, the reduction is straightforward.
Given an instance of MkU, we define a corresponding instance of the Adversarial Missingness for Minimizing the Bias in the Data (AMMBD) problem as follows. We represent each item in $\mathcal{I}$ and each set in $\mathcal{S}$ as a data point $x$ and assign all data points to both the non-sensitive group ($s_{v}=0$) and the positive class ($y=1$). Moreover, we will add one data point $x^{\prime}$ that will belong to the sensitive group ($s_{v}=1$) and the positive class ($y=1$). Therefore, the resulting set of data points is $\{x_{1},x_{2},\ldots x_{n}\}$, where $n=|\mathcal{S}|+|\mathcal{I}|+1$. Moreover, let $x_{i}$ cover $x_{j}$ iff $x_{i}$ represents a set $S_{i}\in\mathcal{S}$, $x_{j}$ represents an item $i_{j}\in\mathcal{I}$, and $i_{j}\in S_{i}$. It follows that minimizing the bias in the AMMBD instance by selecting data points $x_{j}$ to be labeled is equivalent to minimizing the coverage in the corresponding MkU instance.
In terms of approximation, the best-known approximation algorithm for the minimum k-union problem used in our reduction is $O(n^{1/4+\epsilon})$, where $n$ is the number of sets and $\epsilon>0.$, which is mapped to the size of the graph in our problem [13]. This means that at least $O(kn^{1/4})$ nodes would have to be selected to achieve the optimal coverage for $k$ nodes. We note that this is a worse approximation than the classical $O(\log(n))$-approximation for set cover achieved by a greedy algorithm.
Theorem 2
Proof.
From Eq. 7, the min-max objective of BFtS is:
$\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}$ $[\log(f_{A}(\mathbf{h}))]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))]$
$=\sum_{h}p(\mathbf{h}|\hat{s}=1)\log(f_{A}(\mathbf{h}))+\sum_{h}p(\mathbf{h}|\hat{s}=0)\log(1-f_{A}(\mathbf{h}))$
If we fix $\theta_{C}$ and $\theta_{I}$, then the optimal $f_{A}^{*}$ is $\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}.$ By substituting the optimal $f_{A}^{*}$ in Eq. 7 we get:
$\min_{\theta_{C}}\max_{\theta_{I}}$~ $\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}]+$
$\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)})]$
We further simplify the objective function and get:
$\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}$ $[\log\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}]+$
$\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log\frac{p(\mathbf{h}|\hat{s}=0)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}]$
$=\sum_{h}p(\mathbf{h}|$ $\hat{s}=1)\log\frac{p(\mathbf{h}|\hat{s}=1)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}+$
$\sum_{h}p(\mathbf{h}|\hat{s}=0)\log\frac{p(\mathbf{h}|\hat{s}=0)}{p(\mathbf{h}|\hat{s}=1)+p(\mathbf{h}|\hat{s}=0)}$
$=-\log 4$ $+2JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$
By removing the constants, we can further simplify the min-max optimization with optimal $f_{A}^{*}$ to:
$\min_{\theta_{C}}\max_{\theta_{I}}JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$
The optimal $f_{I}^{*}$ thereby maximizes the following objective:
$JS(p(\mathbf{h}|\hat{s}=1);p(\mathbf{h}|\hat{s}=0))$
Since $\hat{y}=\sigma(\mathbf{h.w})$, $f^{*}_{I}$ generates imputation so that
$\min_{\theta_{C}}\sup_{\hat{s}}JS(p(\hat{y}|\hat{s}=1);p(\hat{y}|\hat{s}=0))$
The supremum value will make the two probabilities maximally different. Therefore, for the optimal $f^{*}_{I}$ we get,
$\min_{\theta_{C}}\sup_{\hat{s}}|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$
Let $\Delta DP_{f_{C},f_{I}}=|p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|\hat{s}=0)|$, then, the optimal $f^{*}_{I}$ generates the worst-case (minimum fairness) imputation so that $\Delta DP_{f_{C}.f^{*}_{I}}\geq\Delta DP_{f_{C},f^{*}_{I}}$
Theorem 3
Proof.
From Theorem 1, we get,
$\min_{\theta_{C}}\sup_{\hat{s}}|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$
Let us assume that there can be $m$ different $\hat{s}$ that produce sensitive attributes $\hat{s}_{1},\hat{s}_{2},...,\hat{s}_{m}$. If we let $D_{i}$ be the uniform distribution over each $s_{i}$, then BFtS minimizes the $|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$ in the worst case over these different distributions $D_{i}$. Therefore, $f^{*}_{C}$ is a minimax estimator. The objective of Eq. 7 with optimal $f^{*}A,f^{*}I$ and $f^{*}_{C}$ becomes:
$\inf_{\hat{y}}\sup_{\hat{s}}|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$
Therefore, $f_{C}$ is a minimax estimator. As $f_{C}$ is a minimax estimator, the maximal $\Delta DP$ of BFtS is minimum amongst all estimators of $s$
At the global minima of $f_{C}$, $|p(\hat{y}|\hat{s}=1)-p(\hat{y}|\hat{s}=0)|$ will be minimum. Its minimum value is 0. Therefore, $p(\mathbf{h}|\hat{s}=1)-p(\mathbf{h}|\hat{s}=0)=0$. The optimal $f^{*}_{C}$ achieves demographic parity for the worst case.
Corollary 1
Proof.
If $s^{\prime}$ is a separate imputation independent of $f_{C}$ and $f_{A}$, then the min max game according to [14] is the following:
$\begin{split}\min_{\theta_{C}}\max_{\theta_{A}}~&\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|s^{\prime}=1)}[\log f_{A}(\mathbf{h})]\\ &+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|s^{\prime}=0)}[\log(1-f_{A}(\mathbf{h}))].\end{split}$
With optimal $f^{*}_{A}$, the optimization problem simplifies to-
$\min_{\theta_{C}}-\log 4+JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0)).$
With independent imputation $s^{\prime}$, the adversary $f_{A}$ tries to approximate the lower bound of Jensen Shannon Divergence [71]. In BFtS, the adversarial imputation $f_{I}$ tries to approximate the upper bound of the JS divergence, and the classifier $f_{C}$ tries to minimize the upper bound on the JS divergence provided by the adversarial imputation (see Proposition 1). Therefore,
$JS((p(\mathbf{h}|s^{\prime}\!=\!1);p(\mathbf{h}|s^{\prime}\!=\!0)))\!\leq\!JS(p(\mathbf{h}|\hat{s}\!=\!1);p(\mathbf{h}|\hat{s}\!=\!0)).$
Convergence of adversarial learning for three players vs two players with independent imputation
Let us consider a scenario when the independent missing sensitive value imputation [14] fails to converge. There are three possible reasons for the convergence to fail:
  1. 1.
    $s^{\prime}$ is always 11
  2. 2.
    $s^{\prime}$ is always 0
  3. 3.
    $s^{\prime}$ is uniform, i.e. $p(\mathbf{h}|s^{\prime}=1)=p(\mathbf{h}|s^{\prime}=0)$
In scenario $1$ and $2$ the supports of $p(\mathbf{h}|s^{\prime}=1)$ and $p(\mathbf{h}|s^{\prime}=0)$ are disjoint. According to [14], the optimization for the classifier with an optimal adversary is
$\min_{\theta_{C}}-\log 4+2JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$
As the supports of $p(\mathbf{h}|s^{\prime}=1)$ and $p(\mathbf{h}|s^{\prime}=0)$ are disjoint, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is always 0. The gradient of the JS divergence vanishes, and the classifier gets no useful gradient information—it will minimize a constant function. This results in extremely slow training of the node classifier, and it may not converge.
In scenario 3, $\hat{s}$ is uniform. Therefore, $p(\mathbf{h}|s^{\prime}=1)$ will be equal to $p(\mathbf{h}|s^{\prime}=0)$ which results in $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ being equal to 0. Moreover, the classifier may assign all training samples to a single class and yet have an adversary that may not be able to distinguish the sensitive attributes of the samples as $\hat{s}$ is random. If the classifier assigns all samples to the majority class, it will achieve low classification loss $L_{C}$ along with minimum $L_{A}$. The training of the classifier may get stuck in this local minima (see Eq. 4). This phenomenon is similar to the mode collapse of GANs [65].
Based on Corollary 1, our approach is more robust to the convergence issues described above.
Training Algorithm